Required 1. Determine the dollar amount of DimLoks present annual fixed costs per year. 2. Determine the number of units thaRequired information [The following information applies to the questions displayed belowj RayLok Incorporated has invented aComplete this question by entering your answers in the tabs below. Req 1 Req 2 Req 3a Req 3b Determine the dollar amount of DComplete this question by entering your answers in the tabs belovw Req 1 Req 2 Req 3a Req 3b Determine the number of units thComplete this question by entering your answers in the tabs Req 1 Req 2 Req 3a Req 3b Without regard to your answer in requir
Answer #1

Answer Page No 0 Demkoks present annual tested events porn your 0 Dollar amount of Dollar amount of Dimhors present annualPage, wo @ No. of units = fixed costs + desised profit Contribution margir per unit :4 200,000 + 840,000 $210 - (100+30) 5040Page No ③ Contribution onits gange of 44000 to 64000 units Contribution margir per unit = 180 - [160+30) = 180-130: = $50 No.- Page No 4 Variable coof per unit for coils in excess og 64000 coils, Contribution mango per unit a $180 - (#100 - 30+30) =budached income talement @@ The budgeted produce alderent - Page No ③ Revenue = (44000 x $ 210) + ( 4 2100 $ 180) = 9240000 +Page No ① Total clesived profit = (aoy x fized cont) + Rrofit per unit x Hofal emits sold) = (204.5240,000) + $30 x [64000 42